+1 (585) 613-2221:

with a population size of 50 i want to test two claims on two differenf variables. the 1st claim is to test if the mean of my sample will be less than or equal to 300. the second calim is the see if the mean is equal to 25. then test them and see if we reject the null or fail to reject.

mean1= 279.99

mean2= 20.1

sd1= 321.88

sd2= 22.08
Answer #1

Answer: (i) 18t claim Crven, mean = 279.99 Standard deveation = 321.88 sample size LA) = 50. The mean of my sample will be orP-value: C Pralue for the test statistic is the standard P-value=P(2= -0.44) From the nomal table = 0.32997 ~ 0, 3300 IP-valudih and charm Given, mean (a) = 20.11 Standard dewatson (o)- 22. of samplesrein) = 50. Mean or equal to 25 Wull and altanativP-value; P-value for the test stads stiefs P-value = P(2= -1.57) v From the standard nolmal table. 2 0.05821 ~ 0.0582 [P-valu